OpenStudy (anonymous):

In a scale drawing of Samantha's apartment, the ratio is 1 in = 2 ft. The length of the living room on the drawing is 10 in and the width is 2 in. What is the perimeter of Samantha's living room?

OpenStudy (anonymous):

A 48 ft B 24 ft C 36 ft

jimthompson5910 (jim_thompson5910):

1/2 = 10/L solve for L and tell me what you get

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

1*L = 2*10 L = ???

OpenStudy (anonymous):

20

jimthompson5910 (jim_thompson5910):

1/2 = 2/W W = ??

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

add them up and double the result 20+4 = 24 2*24 = 48

jimthompson5910 (jim_thompson5910):

so the perimeter is 48